### v2

# AgenticIE: An Adaptive Agent for Information Extraction from Complex Regulatory Documents

Abstract
Declaration of Performance (DoP) documents, mandated by EU regulation, certify the performance of construction products. There are two challenges to make DoPs machine and human accessible through automated key-value pair extraction (KVP) and question answering (QA): (1) While some of their content is standardized, DoPs vary widely in layout, schema, and format; (2) Both users and documents are multilingual. Existing static or LLM-only Information Extraction (IE) pipelines fail to adapt to this structural document and user diversity. Our domain-specific, agentic system addresses these challenges through a planner-executor-responder architecture. The system infers user intent, detects document language and modality, and orchestrates tools dynamically for robust, traceable reasoning while avoiding tool misuse or execution loops. Our agent outperforms baselines (ROUGE: 0.783 vs. 0.703/0.608) with better cross-lingual stability (17-point vs. 21-26-point variation).
envname-P envname#1
AgenticIE: An Adaptive Agent for Information Extraction from Complex Regulatory Documents
Gaye Colakoglu1,2 Gürkan Solmaz2 Jonathan Fürst1 1Zurich University of Applied Sciences, Switzerland 2NEC Laboratories Europe, Heidelberg, Germany colgay01@students.zhaw.ch, guerkan.solmaz@neclab.eu, jonathan.fuerst@zhaw.ch
1 Introduction
Declaration of Performance (DoP) documents, mandated under the EU Construction Products Regulation (CPR) European Commission (2014), certify that construction products and materials meet defined performance criteria. For example, these documents often specify the fire protection class. This information is used by the building engineers and architects to select appropriate materials that fit to the specific context (geographical location, local legislation, in-building placement…). Though EU regulation specifies required content it does not define a specific data model and format: DoPs may be scanned or digital, structured or semi-structured, and multilingual (e.g., German or English). Layouts differ by manufacturer, with key information dispersed across tables, free-form text, and hybrid formats.
Automating KVP extraction and QA from DoPs helps reduce manual work, improve digital compliance processes and enable an automated quality control in building construction (e.g., through Building Information Modeling, BIM Garrido-Hidalgo et al. (2022)). But existing pipelines often struggle with noisy scans, inconsistent layouts, and multiple languages Xu et al. (2020); Kim et al. (2021); Sarkhel and Nandi (2021); Perot et al. (2023). Unlike single-shot text- or vision-only systems, agentic systems can track decisions, recover from errors, and adapt tools dynamically Nooralahzadeh et al. (2024). Our intuition with this work is that these characteristics make them a better fit for IE tasks from complex, real-world documents.
Towards this goal, we propose a domain-specific agentic system for DoP documents that infers user intent (KVP or QA), detects modality (scanned or digital), and selects tools via a planner-executor-responder loop( 2). Inspired by tool-augmented agents Schick et al. (2023), modular planners Liu et al. (2023b), and structured frameworks Wu et al. (2024), it maintains state and adapts to user goals and document structure for improved robustness. The agent outperforms baselines across nearly all configurations, achieving ROUGE scores of 0.783 vs. 0.703 (GPT-4o, +11%) and 0.608 (GPT-4o-V, +29%). Further, we release a manually annotated DoP dataset with human-labeled KVP and QA annotations. Experiments show our agent outperforms GPT-4o and vision-based pipelines, especially on diverse layouts and languages.
In summary, our main contributions are:
We release the first, multilingual (english, german) dataset of 80 DoP documents with overall more than 10K human-labeled KVP and QA annotations (Section LABEL:sec:data_construction).
We design an agentic system that adapts to user intent and document modality for Information Extraction for KVP and QA tasks on DoP documents (Section LABEL:sec:agentic_dop_understanding).
We extensively evaluate KVP and QA against GPT-4o text and vision baselines (format validity, schema validity and value validity), demonstrating improved generalization across layout and language variations (Section LABEL:sec:experimental_setup and 3).
Last, upon acceptance we will share all our code and annotated DoP dataset for usage and reproduction by the research community.
2 Related Work
Visually Rich Document Understanding. VRDU supports tasks such as invoice processing, identity verification, and form parsing. Advances have been driven by layout-aware models Xu et al. (2020, 2021) and multimodal datasets Jaume et al. (2019); Park et al. (2019); Huang et al. (2019) that combine text, layout, and visual features for structured extraction. However, these systems are typically trained on clean, monolingual, and semantically regular documents. In contrast, DoP documents exhibit high variability in layout, quality, and language, making them underrepresented and challenging for standard VRDU pipelines.
Prompt-Based and End-to-End Extraction. Prompt-based methods extend document understanding by framing extraction as question answering Hu et al. (2023), incorporating layout structure Lin et al. (2023), or directly decoding fields from images Kim et al. (2021). While these approaches reduce reliance on rigid schemas, they often assume consistently formatted, high-quality input and have recently shown to require a customization of the extraction pipeline to the specific task and dataset Colakoglu et al. (2025).
Tool-Augmented Reasoning and Modular Agents. Recent work enhances LLMs with tool use for modular and adaptive reasoning Schick et al. (2023); Wang and Duan (2024), but most document pipelines remain static, relying on fixed OCR and parsing sequences Luo et al. (2024); Perot et al. (2023). More flexible designs Zhang et al. (2024a); Neupane et al. (2025) introduce partial modularity (e.g., knowledge guidance or layout-aware prompts) but lack runtime planning or tool orchestration. Agentic systems Patil et al. (2023); Qin et al. (2023); Gao et al. (2023) decouple planning, tool use, and response generation via dynamic API calls or subprograms. While benchmarks such as AgentBench Liu et al. (2023b) evaluate such agents in interactive environments, they do not address structured, document-grounded tasks.
Planning, Feedback, and Memory in Agents. Surveys of LLM agents highlight mechanisms such as task decomposition, plan revision, and tool routing Huang et al. (2024); Masterman et al. (2024). Iterative refinement methods implement feedback loops for recovery Shinn et al. (2023); Madaan et al. (2023); Liu et al. (2023a), which our agent mirrors via a planning-execution-feedback cycle to handle OCR or parsing failures.
Structured planning frameworks support modular reasoning and collaborative workflows Wang et al. (2023); Wu et al. (2024), while code-trained LLMs offer stronger schema-aware planning Yang et al. (2024), motivating our choice of a structured planner. Long-term memory mechanisms-such as scratchpads or persistent state Zhang et al. (2024b)-are integrated in our AgentState to track modality, task intent, and tool outcomes. Retrieval-augmented planners Ruan et al. (2023); Kong et al. (2023) also inform our approach to runtime tool selection across diverse document types.
3 Experimental Evaluation
We evaluate our system in comparison to the baseline methods separately for Key Value Pair (KVP) extraction (Section 3.1) and Question Answering (QA) (Section 3.2). Last we investigate runtime, cost and token usage (Section 3.3).
3.1 Results for the KVP Task
In this section, we evaluate model performance on the KVP extraction using the 12 primary keys defined in Section LABEL:sec:target_fields. Two of these keys contain nested values; that is, they represent additional layers of KVP with varying depth and no fixed structure. We report evaluation results under two complementary settings: fixed-schema and open-schema. The fixed-schema covers all key-value pairs, including nested ones (e.g., valuede in Listing LABEL:lst:kvp_annotation_example), by treating their values as flat strings. The open-schema evaluation, by contrast, focuses only on nested key-value pairs and respects their hierarchical organization when computing performance. A more detailed evaluation of nested structures is provided in Appendix A.3.5.
JSON Validity.
As shown in Table 1, the agentic system demonstrates perfect reliability, consistently producing 100% valid JSON across all language and schema configurations. In contrast, the baseline models show some issues. GPT-4o’s validity slightly drops to 96.4% when processing English documents with a German user query. Interestingly, the vision-based GPT-4o-V struggles more, with its performance degrading noticeably when processing German documents, dropping as low as 94.1%. This is likely due to the added complexity of visually parsing the document and composing structured JSON without direct access to clean text, making it more error-prone in syntax-sensitive tasks.
| GPT-4o | GPT-4o-V | Agent | |||||
| en | de | en | de | en | de | ||
| Fixed Schema | en | 1.000 | 1.000 | 1.000 | 0.961 | 1.000 | 1.000 |
| de | 0.964 | 1.000 | 1.000 | 0.941 | 1.000 | 1.000 | |
| Open Schema | en | 1.000 | 1.000 | 1.000 | 0.962 | 1.000 | 1.000 |
| de | 0.964 | 1.000 | 1.000 | 0.942 | 1.000 | 1.000 | |
Key Match Ratio.
This metric evaluates the structural correctness of model outputs by measuring the overlap between predicted and ground-truth keys. The analysis in Table 2 distinguishes between Fixed Schema keys, where keys are known beforehand and provided in the prompt, and Open Schema keys, where nested keys are not predefined and possibly different for each document.
In the Fixed Schema setting, the agent demonstrates perfect reliability, achieving a Key Match Ratio of 1.000 across all document and query language configurations. The baseline models also perform strongly, although GPT-4o shows a minor degradation in a cross-lingual scenario (0.964), and GPT-4o-V’s performance is slightly reduced when processing German documents (down to 0.941).
The results for the Open Schema task are more revealing. Under these more challenging conditions, the agent’s primary advantage emerges: superior consistency and robustness. While the baseline models’ performance is highly erratic and degrades significantly in cross-lingual settings-with GPT-4o dropping to 0.470 and GPT-4o-V to 0.386-the agent’s scores remain stable within a much tighter range (0.571 to 0.654). This highlights the agent’s resilience, as it substantially outperforms the baselines in challenging mixed-language scenarios. The findings suggest that the agent’s architecture is better suited to handle ambiguity, making it more reliable for tasks where a schema is not explicitly defined.
| GPT-4o | GPT-4o-V | Agent | |||||
| en | de | en | de | en | de | ||
| Fixed Schema | en | 1.000 | 1.000 | 1.000 | 0.961 | 1.000 | 1.000 |
| de | 0.964 | 1.000 | 1.000 | 0.941 | 1.000 | 1.000 | |
| Open Schema | en | 0.596 | 0.470 | 0.640 | 0.386 | 0.592 | 0.633 |
| de | 0.485 | 0.677 | 0.498 | 0.526 | 0.571 | 0.654 | |
Value Similarity.
Finally, we evaluate the semantic accuracy of the extracted values for correctly identified keys, using Exact Match, BLEU, and ROUGE metrics. The results in Table 3 show that the agent consistently outperforms both baseline models across nearly all configurations and metrics.
In the Fixed Schema setting, the agent’s superiority is clear. Its advantage is particularly pronounced in the monolingual German task (, ), where it achieves an Exact Match score of 0.583—a substantial improvement over GPT-4o’s 0.430 and GPT-4o-V’s 0.387. This trend holds for ROUGE scores, where the agent reaches 0.783, again significantly outperforming the baselines. This indicates a more robust capability for accurate value extraction in non-English contexts.
This performance advantage is maintained and often amplified in the more challenging Open Schema task. As this evaluation is conditioned on structural correctness, the agent’s superior Key Match Ratio provides a stronger foundation, leading to better value scores. For instance, in the cross-lingual setting of a German query on an English document, the agent’s ROUGE score of 0.369 is considerably higher than GPT-4o’s 0.207. While BLEU scores are low for all models in this setting—suggesting difficulty with precise phrasing—the agent still consistently achieves the highest scores, indicating a greater degree of textual fidelity in its outputs. Overall, the findings demonstrate that the agent’s architecture excels not only at identifying the correct information structure but also at more accurately extracting its semantic content.
| GPT-4o | GPT-4o-V | Agent | ||||||
| Metric | en | de | en | de | en | de | ||
| Fixed Schema | Exact | en | 0.351 | 0.242 | 0.396 | 0.219 | 0.423 | 0.292 |
| de | 0.310 | 0.430 | 0.307 | 0.387 | 0.346 | 0.583 | ||
| BLEU | en | 0.361 | 0.231 | 0.381 | 0.212 | 0.410 | 0.305 | |
| de | 0.234 | 0.386 | 0.247 | 0.334 | 0.331 | 0.443 | ||
| ROUGE | en | 0.642 | 0.448 | 0.665 | 0.402 | 0.708 | 0.613 | |
| de | 0.504 | 0.703 | 0.504 | 0.608 | 0.666 | 0.783 | ||
| Open Schema | Exact | en | 0.258 | 0.143 | 0.310 | 0.136 | 0.320 | 0.352 |
| de | 0.189 | 0.371 | 0.224 | 0.298 | 0.244 | 0.367 | ||
| BLEU | en | 0.130 | 0.099 | 0.162 | 0.102 | 0.205 | 0.158 | |
| de | 0.091 | 0.089 | 0.121 | 0.183 | 0.195 | 0.189 | ||
| ROUGE | en | 0.307 | 0.467 | 0.381 | 0.347 | 0.387 | 0.452 | |
| de | 0.207 | 0.412 | 0.255 | 0.477 | 0.369 | 0.512 | ||
3.2 Results for the QA Task
We evaluate the QA capabilities using three complementary metrics, as in the KVP evaluation. Table 4 presents comparative performance across models, evaluation scopes, and languages.
| GPT-4o | GPT-4o-V | Agent | ||||||
| Metric | en | de | en | de | en | de | ||
| Fixed Schema | Exact | en | 0.503 | 0.393 | 0.500 | 0.379 | 0.570 | 0.329 |
| de | 0.430 | 0.444 | 0.430 | 0.453 | 0.431 | 0.515 | ||
| BLEU | en | 0.442 | 0.324 | 0.377 | 0.296 | 0.491 | 0.347 | |
| de | 0.313 | 0.374 | 0.313 | 0.382 | 0.401 | 0.487 | ||
| ROUGE | en | 0.726 | 0.568 | 0.674 | 0.519 | 0.780 | 0.655 | |
| de | 0.567 | 0.644 | 0.559 | 0.637 | 0.725 | 0.732 | ||
| Open Schema | Exact | en | 0.367 | 0.292 | 0.262 | 0.212 | 0.384 | 0.384 |
| de | 0.307 | 0.446 | 0.211 | 0.308 | 0.307 | 0.479 | ||
| BLEU | en | 0.244 | 0.178 | 0.160 | 0.125 | 0.254 | 0.245 | |
| de | 0.189 | 0.318 | 0.096 | 0.209 | 0.205 | 0.336 | ||
| ROUGE | en | 0.593 | 0.535 | 0.403 | 0.360 | 0.666 | 0.620 | |
| de | 0.478 | 0.642 | 0.308 | 0.446 | 0.576 | 0.678 | ||
All models perform better when query and document languages align. In aligned scenarios (en/en and de/de), the agent achieves moderate advantages over GPT-4o in fixed schema: +6.7 points Exact Match (0.570 vs 0.503 for en; 0.515 vs 0.444 for de), +5.4-8.8 ROUGE. GPT-4o-V matches GPT-4o when aligned, but degrades more in cross-lingual scenarios.
Cross-lingual asymmetry appears consistently: de/en degrades more than en/de. For the agent in fixed schema, this yields 0.329 vs 0.431 Exact Match (24% relative gap); GPT-4o shows 0.393 vs 0.430 (9% gap), suggesting German document answer extraction is systematically harder with English queries.
In open schema, performance drops substantially. Agent Exact Match spans 0.307-0.479 across conditions; GPT-4o: 0.292-0.446; GPT-4o-V: 0.211-0.308 (uniformly lower). The agent’s advantage over GPT-4o narrows from +6.7 points in fixed schema (aligned en) to +3.3 points in open schema (de/de: 0.479 vs 0.446), indicating validation-based orchestration benefits more from schema guidance.
GPT-4o-V underperforms text approaches in open schema (0.212 Exact Match for de/en vs 0.292 GPT-4o), with gaps widening in cross-lingual settings (0.211 for en/de vs 0.307 GPT-4o), suggesting vision processing compounds nested structure discovery challenges.
3.3 Runtime, Cost and Token Comparison
We further compare the computational efficiency of the agentic system against both baseline models in terms of token usage, total cost, and runtime. Table 5 summarizes these metrics across the KVP and QA tasks. Input and output token counts are derived from the prompt and completion segments, respectively, with visual or textual context included in input estimates. The cost is computed based on the GPT-4o pricing (Sept. 2025), and the run time reflects the total wall clock time per tasks. Although the agentic system provides performance benefits (see Section 3.1 and 3.2), the cost in terms of runtime and token usage is higher, which is an open research problem Nooralahzadeh et al. (2024) and needs to be solved for sustainable agentic IE.
| GPT-4o | GPT-4o-V | Agent | ||
| Total Tokens (M) | 3.21 | 21.22 | 38.69 | |
| Total Cost (USD) | 18.15 | 195.88 | 312.59 | |
| Runtime (min) | 89.8 | 430.7 | 1180.2 |
4 Conclusion
This paper introduces an agentic system for IE from DoP documents, with a focus on handling hierarchical structures in English and German. Our agent consistently generates 100% valid JSON and outperforms GPT-4 baselines in matching nested keys. The agent’s primary strength is nested extraction stability, with performance varying less than 9% across configurations, whereas baselines fluctuate by up to 25%.
Limitations
This work focuses on adaptive information extraction within a specific regulatory context-Declaration of Performance (DoP) documents-characterized by multilinguality and diverse layouts. The agentic system is therefore intentionally designed for domain-grounded reasoning rather than general-purpose intelligence. Extending it to other domains such as financial or legal records would require retraining the planner and adapting the tool registry to new schemas and reasoning conventions.
Our evaluation emphasizes structural and semantic correctness but does not address higher-order reasoning such as factual consistency, regulatory validation, or human-in-the-loop correction. Exploring such capabilities would shift the system from extraction toward auditable compliance reasoning. Likewise, while the agent performs reliably across English and German, scaling to low-resource or morphologically complex languages would demand cross-lingual schema alignment and translation-aware planning.
Finally, the current design focuses on textual content and inferred structure. Incorporating explicit visual cues such as layout topology or table geometry could further enhance precision in complex or scanned documents, but doing so would require multimodal model integration beyond this paper’s technical scope.
Ethical Considerations
The development and deployment of this agentic system for extracting information from Declaration of Performance (DoP) documents require careful consideration of several ethical dimensions, including data privacy, potential for misuse, algorithmic bias, and environmental impact.
Data Privacy and Personal Information.
As noted, DoP documents are public records mandated by EU regulation, but they can contain personal identifiers, such as the names, positions, and signatures of the individuals responsible for the declaration. While this information is publicly accessible, its automated collection and processing at scale warrant ethical scrutiny. Our approach mitigates privacy risks by using this data strictly for research purposes to train and evaluate the extraction system. We have not released any raw documents or personally identifiable information (PII). Furthermore, all annotations were performed by the authors, involving no external human subjects. To maintain this standard, any future release of the dataset will be carefully reviewed to ensure compliance with privacy best practices.
Reliability and Risk of Misuse
The information extracted from DoPs, such as fire resistance ratings or the use of hazardous substances, is critical for ensuring the safety and compliance of construction projects. An error in extraction could lead to the selection of non-compliant materials, posing significant safety risks. Therefore, this system is intended to be an assistive tool for engineers and architects, not a fully autonomous replacement for human oversight. End-users should be made aware that the system’s output must be verified by a qualified professional before being used for any critical decisions. The system’s built-in verification steps and high performance are designed to minimize errors, but the "human-in-the-loop" remains essential for accountability in high-stakes applications.
Algorithmic Bias and Fairness
Our dataset comprises documents in English and German. While the agentic system demonstrates more stable cross-lingual performance than the baselines, the paper notes that performance can still vary between languages and in cross-lingual scenarios. Deploying such a system across the multilingual European market could lead to performance disparities, potentially disadvantaging users working with less-represented languages. The data was also collected via web search and from specific repositories, which may introduce a selection bias towards certain manufacturers or product types. Future work should aim to expand the dataset to include a wider range of languages and document origins to ensure more equitable and robust performance.
Computational Cost and Environmental Impact
The experiments show that the proposed agentic system, while more accurate, incurs a significantly higher computational cost in terms of token usage, cost, and runtime compared to baseline models. The total token count for the agent was 38.69 million, costing over $300 and taking nearly 20 hours to run. We acknowledge that the energy consumption associated with such extensive computation has an environmental impact. This trade-off between performance and efficiency is a known challenge for agentic architectures. We advocate for continued research into more computationally efficient models and methods to create more sustainable solutions for information extraction.
AI Assistants In Research Or Writing
We have used GPT-5, Gemini 2.5 pro and Claude Opus 4.1 to help us refine our writing. All final text in the paper has been verified by the authors.
Appendix A Appendix
A.1 Dataset Annotation Details
A.1.1 Target Keys
The following JSON-style structure presents list of 12 target keys selected for annotation and extraction from DoP documents, including their English and German versions. These keys represent core elements required for regulatory compliance.
A.1.2 KVP Annotation Sample
Example of a multilingual KVP annotation for the key "Declared Performance" from a German-language DoP document.
The following JSON-style structure presents extracted key-value pairs (KVPs) from the document Sager_ISO-SWISS-Design-Panels.pdf, annotated in both English and German for evaluation and training purposes.
A.1.3 QA Annotation Sample
The following JSON-style structure presents a sample of question annotations derived from key-value pairs in the document 011-Leistungserklaerung-KS-10DF.pdf. Each entry specifies the key, its value, the associated question, and the language used.
A.2 Agentic System Implementation Details
A.2.1 AgentState Variables
A.2.2 AgentStatus Variables
| AgentStatus | Description |
| PLAN | Reason about the next step |
| NEED_TOOL | A tool has been selected; execute it |
| RESPOND | Enough information gathered; generate final output |
| SUCCESS | Final answer produced successfully |
| END | Workflow terminated (successfully or forcibly) |
A.2.3 AgentState and Execution Flow Diagram
A.2.4 Tool Registry and Modularization
To support robust and adaptable behavior across multilingual, layout-diverse, and partially scanned DoP documents, system relies on a modular tool architecture. Tools in this system represent discrete capabilities, such as PDF modality detection, OCR-based text extraction, question answering, or key-value pair retrieval, and are invoked selectively based on the agent’s current reasoning state.
At the heart of this architecture is a structured Tool abstraction, which wraps both deterministic utilities (e.g., OCR, PDF parsing) and LLM-powered capabilities (e.g., intent classification or verification) into a unified callable interface. Each tool object declares its name, functional implementation, Pydantic input schema, and optional pre- and post-processing logic. Tools also specify whether they require previous output, and which document types they are compatible with (e.g., scanned or text-based). This explicit declaration ensures that tools remain self-contained and declarative, while still integrating seamlessly into the planner-executor loop.
All tools are registered at startup by combining deterministic and LLM-based implementations into a central registry, forming a complete vocabulary of system capabilities. This registry remains fixed throughout execution and provides a filtered view of what actions are currently applicable, based on the evolving AgentState.
When the planner is invoked, it consults this registry to determine which tools are compatible with the current reasoning intent and document modality. Tools that require unavailable inputs (e.g., translation before text extraction) are excluded. The remaining candidates are embedded as tool metadata into the LLM prompt. The model then selects the most appropriate tool and input parameters via structured output, which is parsed and validated against the registry before dispatching it to the executor.
The executor interprets this plan by invoking the selected tool. Before execution, it applies a normalize_input() function to map the current state and raw parameters into a validated schema. Once executed, the result is post-processed via the tool’s postprocess() function, which updates the AgentState in a localized and semantically meaningful way. This decoupling of decision logic from execution logic ensures that tools remain modular, reusable, and interchangeable, while still preserving contextual awareness.
Unlike monolithic pipelines, the tools in the system are deeply integrated into a loop of stateful reasoning. For instance, modality detection tools such as check_if_scanned determine whether a document should be OCR’d or parsed directly, triggering either extract_text_ocr or extract_text_direct accordingly. Language inference via detect_language enables downstream translation and schema matching. Intent classification and target key parsing refine the agent’s task trajectory, while extraction tools like extract_key_values or answer_question carry out the core retrieval logic. The planner may later invoke verify_extraction to assess whether the output is complete and grounded in the source text, before optionally translating it into the user’s preferred language using translate_text. In edge cases, where noisy OCR output risks triggering LLM content filters, the system sanitizes text using sanitize_ocr_text as a recovery mechanism.
Each of these tools serves a specific role in a broader adaptive plan. Their activation is not hardcoded but emerges dynamically as the agent iterates through reasoning, acting, and revising. Tools modify the state incrementally, and their execution history is logged to support loop detection, misuse prevention, and multi-step recovery. Because each tool operates independently but updates a shared AgentState, the planner and executor can reason about partial progress, fallback strategies, or when to terminate altogether.
The tight integration of tools, registry, and reasoning enables the agent to dynamically adapt its behavior across a wide variety of document structures and user inputs, while preserving traceability and control. Figure 4 illustrates the complete tool orchestration pipeline, from initial tool registration to planner-driven selection and executor-based integration.
Multilingual QA and KVP Templates: To support multilingual document processing, the system employs language-specific prompt templates for both question answering and key-value extraction tasks. Based on the detected document_language, the agent selects between English (en) and German (de) versions of the prompts. These templates ensure that language constraints, extraction rules, and formatting guidelines are consistently enforced.
The examples below illustrate two representative prompts used during tool invocation: one for KVP extraction in English, and one for QA in German.
Each prompt includes a schema definition and task-specific extraction rules tailored to the document’s language. This modular design ensures that the model generates structured outputs consistent with downstream processing requirements such as verification and translation.
A.2.5 Prompt Design and Language Adaptation
The reasoning capability of our agentic system depends heavily on well-structured and context-aware prompts. These prompts guide both the planner and tool-specific language models to make decisions, extract structured information, and adapt across languages. This section outlines the design principles and implementation of prompt templates for planning, key-value extraction, question answering, error recovery, and multilingual handling.
Planner Prompt with State-Aware Reasoning: The planner prompt is dynamically constructed based on the current AgentState and tool metadata. The main structure is as follows:
It synthesizes the current AgentState into a rich context string that includes user intent, document metadata, extracted content, prior tool usage, tool descriptions, and reasoning history. This state-driven prompt is used to instruct a language model to determine the next action in the pipeline. The planner must always respond in a strict JSON format:
The prompt encodes decision-making rules directly, allowing for multi-step planning logic. For example, if the user intends to extract key-value pairs but the document is in German while the user expects English output, the planner prompt will contain structured rules for translation, extraction, verification, and result transformation:
Consider the state where the user has submitted a document titled DoP-1234.pdf with the intent of extracting key-values. The document has already been classified as scanned, and its language is detected as de (German), while the user’s preferred language is en (English). No extracted values or translated keys exist yet. The planner prompt will reflect these observations as extracted facts:
The planner might then respond with:
This strict format allows the system to introspect, log, and act deterministically on LLM outputs.
Error Recovery Prompts: To handle tool failures or malformed outputs, recovery prompts provide fallbacks. For instance, if the planner returns invalid JSON, the system uses get_malformed_response_prompt():
This prompt is structurally identical to the planner prompt but framed in an error-recovery mode, helping the agent self-correct without external intervention.
By combining language-specific prompt templates, strict JSON response schemas, and a planning prompt with embedded decision rules, the system achieves structured reasoning across multilingual documents. The modularity of this design also enables the easy addition of new languages, extraction schemas, or prompt variations, supporting scalability in diverse information extraction scenarios.
A.2.6 KVP Schema templates for EN/DE
A.3 Agentic System Evaluation Related Details
A.3.1 Metrics Details
Exact Match (EM). provides the strictest evaluation criterion by requiring complete string equality between the predicted value and ground truth :
This metric is particularly important for structured fields such as declaration numbers, product codes, where even a single character deviation renders the prediction unusable. Many such fields in DoP documents follow rigid templates and should be reproduced verbatim.
BLEU Score (with smoothing). measures the -gram precision between the predicted value and the reference . It is effective for short to medium-length technical descriptions. Though originally designed for machine translation, it rewards overlapping lexical fragments even if their order varies, we compute it as:
where is the modified -gram precision for order , is the weight assigned to each -gram (typically ), and BP denotes the brevity penalty to penalize short predictions. BLEU is computed using NLTK’s sentence-level scorer with up to 4-grams and SmoothingFunction (method 4) to handle cases with low or zero overlap.
BLEU is helpful in evaluating fields like intended use or assessment systems, where phrasing may differ while retaining technical terminology. However, it is limited in recall and less robust to paraphrasing, which is addressed by ROUGE-L.
ROUGE-L. measures the similarity between predicted and reference text using the longest common subsequence (LCS). The F1 version of the ROUGE-L score is defined as:
where is the precision, is the recall, and balances precision and recall.
We implement ROUGE-L using the rouge-score Python library with stemming enabled. Unlike BLEU, ROUGE-L emphasizes sequence alignment and recall, making it suitable for long, nested fields such as the "Declared Performance" block. It is robust to minor rewordings, reordering, and partial omissions.
Each metric reveals different correctness and failure modes. Exact Match enforces binary correctness, essential for identifiers. BLEU captures surface-level lexical overlap in moderately variable phrases. ROUGE-L handles paraphrased, nested, or reordered content in longer fields. For example, a prediction might fail EM due to formatting but still convey equivalent meaning, reflected by a high ROUGE score. BLEU helps distinguish semantically plausible but imprecise outputs from incomplete ones.
As shown in Figure 5, the combined use of EM, BLEU, and ROUGE-L ensures robust coverage across content types, ranging from atomic identifiers to free-text paragraphs, balancing strict correctness with semantic flexibility.
After computing JSON Validity, Key Match Ratio (see Equations (LABEL:eq:json), (LABEL:eq:keymatch)), and Value Similarity scores (Exact Match, BLEU, and ROUGE-L (see Equations (1), (2), (3))), we aggregate results at the document and language levels to enable fair comparison across multilingual DoP documents.
A.3.2 Aggregation of Evaluation Metrics for KVP Task
As described in Algorithm 1, evaluation is conducted on each pair, producing five scalar scores. The system is expected to predict values for 12 target keys per document. To ensure this expectation is met and the evaluation is not biased by incomplete predictions, we assign a similarity score of zero for any ground truth key that is missing from the system output. This effectively penalizes omissions. Value similarity metrics (Exact Match, BLEU, ROUGE-L) are computed only on matched keys, and the totals are normalized by to ensure consistency across varying numbers of matches. After per-sample evaluation, we aggregate results by computing document–language averages, followed by a final macro-average over all documents and languages.
A.3.3 Aggregation of Evaluation Metrics for QA Task
For the QA task, the system is evaluated on its ability to correctly answer a set of natural language questions for each pair, as outlined in Algorithm 2. Similar to the KVP task, if a ground truth question exists but the system provides no answer, we assign a score of zero for that item across all metrics. This approach avoids discarding samples with missing predictions and ensures that models are fairly penalized for incomplete outputs. Unlike the KVP task, the number of questions may vary per document (see Section LABEL:sec:qa_data_generation), so no fixed normalization is applied. Instead, we compute unweighted averages of the value similarity scores across all available question–answer pairs for each document–language combination. These document-level scores are then aggregated to obtain overall averages across the evaluation set.
A.3.4 Error Logging and Invalid Outputs
To ensure robustness and traceability in evaluation, all execution failures encountered during agent processing are systematically recorded. These errors may stem from input parsing issues, image extraction failures, model output malformation, or other runtime exceptions. By logging failure metadata such as document name, user language, and error trace, we enable targeted debugging and facilitate quantitative analysis of system limitations.
KVP Task. In the key–value extraction task, errors typically occur at the document level, corrupted PDFs, failed extraction pipelines, or outputs that cannot be parsed into valid JSON structures. Since the agent is required to return a complete JSON structure containing 12 predefined key–value pairs, partial evaluation is not feasible. Even if the process fails midway, the entire output is considered invalid. Rather than discarding these cases from evaluation, we assign zero scores to all metrics. This zero-padding approach ensures consistent treatment across the dataset and prevents artificially inflated averages. To support post-hoc analysis, all failed cases are logged with document metadata and error traces. We also store the raw outputs with validity flags to facilitate deeper investigation of structural errors.
QA Task. For the QA task, each document may contain multiple questions in a given language, and errors are logged at the level of individual question–document–language triples. Failures can arise during PDF-to-image conversion, multi-image model inference, or malformed model outputs (e.g., empty strings or irrelevant text). Unlike KVP, partial evaluation is still possible: if some questions fail, the remaining ones from the same document can still be evaluated. This enables broader coverage even in the presence of localized failures.
In both tasks, structured error logging plays a critical role in exposing systemic weaknesses and guiding future improvements to the agent’s robustness and reliability across languages and document formats.
A.3.5 Aggregation of Evaluation Metrics for Nested-KVP Task
The evaluation metrics introduced so far-JSON Validity, Key Match Ratio, and Value Similarity scores such as Exact Match, BLEU, and ROUGE-L ( (2), (1), (LABEL:eq:keymatch), (LABEL:eq:json), (3))-assume a flat structure, where each target corresponds to a single key–value pair. However, in the KVP task, some fields, such as Declared Performance, Signature, and their German counterparts Erklärte Leistung and Unterschrift, contain structured values encoded as nested JSON objects.
To more accurately evaluate such cases, we extend our framework with a recursive strategy tailored for nested values. As detailed in Algorithm 3, the process begins by verifying that the predicted and ground truth values for these keys are valid JSON objects. If valid, each structure is flattened into a set of path–value pairs, where each path is constructed by joining nested keys with a delimiter (e.g., "/"), as shown below.
This flattening enables consistent alignment and scoring using the same evaluation logic as for flat fields. Metrics are computed for each matched key path using the same formulas: Exact Match, BLEU, and ROUGE-L. These scores are then aggregated at both the document and language levels following the same procedure and the same error logging strategy as in the KVP task.
A.4 Baseline Prompt Examples
A.4.1 KVP Baseline Prompt (Version 1 - English)
A.4.2 KVP Baseline Prompt (Version 2 - German)
A.4.3 QA Baseline Vision Prompt (English)
A.5 Further Result Discussion
| Ground Truth Key (GT) | Predicted Key | Explanation |
| Erklärte Leistung/Klasse der Brutto-Trockenrohdichte | Erklärte Leistung/Brutto-Trockenrohdichte | Qualifier “Klasse der” is missing in the prediction; keys fail to align. |
| Erklärte Leistung/Form und Ausbildung/Bezeichnung | Erklärte Leistung/Form und Ausbildung | Predicted key refers to a parent field; GT specifies a subfield. |
| Erklärte Leistung/Wasserdampfdurchlässigkeit (Tabellenwert) | Erklärte Leistung/Wasserdampfdurchlässigkeit | Suffix “(Tabellenwert)” is omitted; results in an exact match failure. |
| Erklärte Leistung/Maße/Länge [mm] | Erklärte Leistung/Maße | Prediction aggregates multiple subfields; GT expects specific ones. |
| Erklärte Leistung/mittlere Druckfestigkeit [N/mm²] | Erklärte Leistung/mittlere Druckfestigkeit | Unit annotation is missing in the prediction; leads to mismatch. |
| Unterschrift/Ort und Datum der Ausstellung | Unterschrift/Ort und Datum der Ausstellung | Keys match, but predicted value has a spelling error (“Lobnitz” vs. “Löbnitz”). |
| Erklärte Leistung/Wärmedurchlasswiderstand | — | Key missing in prediction; leads to evaluation exclusion. |
| Erklärte Leistung/Form und Ausbildung/Lochanteil von - bis [Vol %] | — | GT key has no corresponding prediction; reduces recall. |
| Erklärte Leistung/Grenzabmaße / Abmaßklasse | — | Missing predicted key causes evaluation penalty. |