### v3

# AgenticIE: An Adaptive Agent for Information Extraction from Complex Regulatory Documents

Abstract
Declaration of Performance (DoP) documents, mandated by EU regulation, specify characteristics of construction products, such as fire resistance and insulation. While this information is essential for quality control and reducing carbon footprints, it is not easily machine readable. Despite content requirements, DoPs exhibit significant variation in layout, schema, and format, further complicated by their multilingual nature. In this work, we propose DoP Key Information Extraction (KIE) and Question Answering (QA) as new NLP challenges. To address this challenge, we design a domain-specific AgenticIE system based on a planner-executor-corresponder pattern. For evaluation, we introduce a high-density, expert-annotated dataset of complex, multi-page regulatory documents in English and German. Unlike standard IE datasets (e.g., FUNSD, CORD) with sparse annotations, our dataset contains over 15K annotated entities, averaging over 190 annotations per document. Our agentic system outperforms static and multimodal LLM baselines, achieving Exact Match (EM) scores of 0.396 vs. 0.342 (GPT-4o, +16%) and 0.314 (GPT-4o-V, +26%) across the KIE and QA tasks. Our experimental analysis validates the benefits of the agentic system, as well as the challenging nature of our new DoP dataset.
envname-P envname#1
AgenticIE: An Adaptive Agent for Information Extraction from Complex Regulatory Documents
Gaye Colakoglu1,2 Gürkan Solmaz2 Jonathan Fürst1 1Zurich University of Applied Sciences, Switzerland 2NEC Laboratories Europe, Heidelberg, Germany colgay01@students.zhaw.ch guerkan.solmaz@neclab.eu jonathan.fuerst@zhaw.ch
1 Introduction
Although EU regulations specify the required content, they do not define a specific data model or format. DoPs can be scanned or digital, structured or semi-structured, and multilingual. Layouts vary by manufacturer, with key information spread across tables, free-form text, and hybrid formats.
To make the contained information in DoP documents machine and human accessible, we propose DoP Information Extraction (IE) as a new application to the NLP community. IE from DoPs reduces manual work, improves digital compliance processes and enables automated quality control in building construction (e.g., through Building Information Modeling, BIM Garrido-Hidalgo et al. (2022)).
Current IE methods for layout-rich documents usually build on custom or general transformer-based models Colakoglu et al. (2025). They can be divided into text-only methods and multi-modal methods. Still, they often struggle with noisy scans, inconsistent layouts, and multiple languages Xu et al. (2020); Kim et al. (2021); Sarkhel and Nandi (2021); Perot et al. (2023). Recently, Agentic AI has been proposed to tackle IE with a diverse document structure LandingAI (2025). Unlike static single-shot text- or vision-only systems, agentic systems can track decisions, recover from errors, and adapt tools dynamically Nooralahzadeh et al. (2024). Our intuition is that these characteristics make them a better fit for IE tasks from complex, real-world documents.
AgenticIE Method. First, we propose a domain-specific agentic system for DoP documents that infers user intent, detects modality (scanned or digital), and selects tools via a planner-executor-responder loop. Inspired by tool-augmented agents Schick et al. (2023), modular planners Liu et al. (2023b), and structured frameworks Wu et al. (2024), it maintains state and adapts to user goals and document structure for improved robustness. To show the working and benefits of AgenticIE, let us consider an example (Figure 1): The user provides an input to the system: “What is the fire resistance rating in this DoP?”; The LLM determines user intent and predicts a QA task; The agent calls a tool to extract the text, which fails immediately as the PDF is image-based; Based on the failed extraction the agent chooses to use an OCR tool to extract the text from the image which succeeds; A language detector tool is called to detect the user and document language which predicts the user question to be in German and the document English; The agent decides to translate the question to English to match the document language; Now with question and document text ready, the QA tool is called to retrieve the fire resistance: 0.01; The answer is verified and returned in user language to the user.
DoP Dataset. Second, we create a new resource of multi-page documents (174 pages) of which we annotate more than 15K key-value and question answer pairs in English and German. Our agentic system outperforms a static LLM and Multimodal LLM baseline, achieving Exact Match (EM) scores of 0.396 vs. 0.342 (GPT-4o, +16%) and 0.314 (GPT-4o-V, +26%) across Key Information Extraction (KIE) and QA tasks.
In summary, our main contributions are:
We release the first, multilingual (English and German) dataset of DoP documents with overall more than 15K human-labeled KIE and QA annotations (Section LABEL:sec:data_construction).
We design an agentic system that adapts to user intent and document modality for KIE and QA tasks on DoP documents (Section LABEL:sec:agentic_dop_understanding).
We extensively evaluate KIE and QA against GPT-4o text and vision baselines (schema and value validity), demonstrating improved generalization across layout and language variations (Section LABEL:sec:experimental_setup and 3).
The AgenticIE code repository and the annotated DoP dataset are available as supplementary files and will be released online for usage and reproduction.
2 DoP IE Problem
IE from DoP documents presents unique difficulties that are distinct from standard visually rich document understanding (VRDU). We identify three key challenges depicted through an example in Figure 2:
[C1] Irregular Table Structure and Cross-References.
DoPs exhibit high layout variation, mixing free-form text with complex, borderless tables spanning multiple pages. A critical challenge is resolving implicit cross-references, where values (e.g., product codes) link to specifications in disjoint tables (Figure 2). Standard OCR models often fail to capture these non-local dependencies, resulting in fragmented extractions.
[C2] Hybrid Schema: Fixed vs. Open Structure.
DoP extraction requires handling hybrid schemas. While standard metadata (e.g., Manufacturer, DoP Number) follow a Fixed Schema where keys are known a priori, technical specifications (e.g., under Declared Performance) follow an Open Schema. In these nested sections, the keys (sub-properties) are dynamic, unknown beforehand, and vary significantly between product types. Systems cannot rely on static templates, but they must dynamically discover and structure these open-ended attributes.
[C3] Multilinguality and Scanned Documents.
DoPs are in different languages, often creating asymmetry, where the user’s query language () differs from the document’s (). Systems must perform on-the-fly semantic alignment to map intents (e.g., English “Fire Resistance”) to native document values (e.g., German “Brandverhalten”).
2.1 Task Definition
We define DoP IE with its two related sub-tasks: Key Information Extraction (KIE) and Question Answering (QA). Let be a document with language and be the user’s target language.
Sub-Task 1: Hybrid KIE
Given and a target schema in , the goal is to extract a structured object . The extraction involves two types of targets: 1. Fixed Keys: Predefined fields (e.g., metadata) provided explicitly to the model. 2. Open Keys: Dynamic fields (e.g., Declared Performance) where the model must autonomously discover a nested substructure without prior knowledge of the sub-keys. The system must preserve this hierarchy and return content in , translating if .
Sub-Task 2: Multilingual QA.
Given and question in , generate answer in . must be grounded in , resolving cross-references and tabular structures.
Precision Requirement.
For both sub-tasks, EM is essential. Unlike general-purpose retrieval where semantic similarity often suffices, the extracted information is used directly to select materials based on specific characteristics (e.g., fire resistance class, thermal conductivity). Consequently, precise extraction of values and units is critical, as approximations could lead to the selection of non-compliant materials.
3 Experimental Evaluation
We evaluate our system in comparison to the baseline methods separately for KIE (Section 3.1) and QA (Section 3.2). Last, we investigate runtime, cost and token usage (Section 3.3).
3.1 Results for the KIE Task
We evaluate extraction performance on the 12 primary keys defined in Section LABEL:sec:target_fields. Consistent with our Hybrid Schema definition, we report results on two distinct subsets of keys: Fixed Schema: The set of standard metadata fields (e.g., Manufacturer, Declaration Number) where the schema is static and known a priori. Open Schema: The complex, nested fields (e.g., Declared Performance) where the agent must dynamically discover the structure and sub-keys without a predefined template.
3.1.1 Schema Adherence.
This metric evaluates structural correctness by measuring the intersection between predicted and ground-truth entity types (Table 1 ).
In the Fixed Schema setting, the Agent achieves perfect reliability (1.0) across all language pairs, effectively solving the metadata extraction task. The baselines are also strong, though GPT-4o-V shows slight degradation (0.941) on German documents, likely due to OCR challenges in the vision encoder.
The Open Schema setting reveals the core advantage of the agentic approach. Discovery of nested keys in Declared Performance is significantly harder, particularly in cross-lingual scenarios. While baselines fluctuate wildly, e.g., GPT-4o drops to 0.470 on En De, the Agent maintains robust performance (0.633), achieving the highest average adherence (0.613). This confirms that the planner-executor loop effectively adapts to dynamic schemas and mixed-language scenarios, where static prompting fails.
| Task | GPT-4o | GPT-4o-V | Agent (Ours) | |
| Fixed Schema | En En | 1.000 | 1.000 | 1.000 |
| En De | 1.000 | 0.961 | 1.000 | |
| De En | 0.964 | 1.000 | 1.000 | |
| De De | 1.000 | 0.941 | 1.000 | |
| Mean | 0.991 | 0.976 | 1.000 | |
| Open Schema | En En | 0.596 | 0.640 | 0.592 |
| En De | 0.470 | 0.386 | 0.633 | |
| De En | 0.485 | 0.498 | 0.571 | |
| De De | 0.677 | 0.526 | 0.654 | |
| Mean | 0.557 | 0.513 | 0.613 |
3.1.2 Exact Match (EM)
We evaluate value correctness using the EM metric. As shown in Table 2, the Agent consistently outperforms baselines, particularly in cross-lingual tasks.
For Fixed Schema keys, the Agent achieves a mean EM of 0.411, surpassing GPT-4o (0.333) and GPT-4o-V (0.327). The gap is most visible in the monolingual German setting (0.583 vs. 0.430), suggesting that the agent’s tool use (specifically tailored OCR and translation tools) yields cleaner value extraction than end-to-end generation.
In the Open Schema task, precision is critical as incorrect nesting invalidates the extracted value. The Agent achieves the highest mean EM (0.321). Notably, in the difficult En De setting (English query, German document), the Agent scores 0.352, more than doubling the performance of GPT-4o (0.143). This demonstrates that our agentic architecture successfully bridges the language gap for complex, nested data structures where baselines struggle to align the user’s English intent with the document’s German hierarchy.
| Task | GPT-4o | GPT-4o-V | Agent (Ours) | |
| Fixed Schema | En En | 0.351 | 0.396 | 0.423 |
| En De | 0.242 | 0.219 | 0.292 | |
| De En | 0.310 | 0.307 | 0.346 | |
| De De | 0.430 | 0.387 | 0.583 | |
| Mean | 0.333 | 0.327 | 0.411 | |
| Open Schema | En En | 0.258 | 0.310 | 0.320 |
| En De | 0.143 | 0.136 | 0.352 | |
| De En | 0.189 | 0.224 | 0.244 | |
| De De | 0.371 | 0.298 | 0.367 | |
| Mean | 0.240 | 0.242 | 0.321 |
3.2 Results for the QA Task
We evaluate QA capabilities using the same EM metric, focusing on the ability to retrieve precise answers for both static metadata (Fixed Schema) and dynamic properties (Open Schema) (Table 3).
| Task | GPT-4o | GPT-4o-V | Agent (Ours) | |
| Fixed Schema | En En | 0.503 | 0.500 | 0.570 |
| En De | 0.393 | 0.379 | 0.329 | |
| De En | 0.430 | 0.430 | 0.431 | |
| De De | 0.444 | 0.453 | 0.515 | |
| Mean | 0.443 | 0.441 | 0.461 | |
| Open Schema | En En | 0.367 | 0.262 | 0.384 |
| En De | 0.292 | 0.212 | 0.384 | |
| De En | 0.307 | 0.211 | 0.307 | |
| De De | 0.446 | 0.308 | 0.479 | |
| Mean | 0.353 | 0.248 | 0.389 |
Fixed Schema QA. Here, language alignment is a strong predictor of performance. For matched languages ( and ), the Agent leads significantly, achieving 0.570 and 0.515 EM scores respectively. This suggests that its explicit verification step filters out hallucinations more effectively than the baselines. A notable exception is the setting (English query on German document), where GPT-4o (0.393) outperforms the Agent (0.329). This dip likely stems from the “translation-then-retrieval” bottleneck: minor errors in translating a specific technical query (e.g., "Notified Body") into German can cause retrieval failures that end-to-end models might bypass through soft alignment.
Open Schema QA. The Open Schema task targets values deeply nested within complex tables (e.g., Declared Performance). Here, the Agent’s architectural benefits are most pronounced. It achieves the highest mean EM score (0.389), outperforming GPT-4o (0.353) and GPT-4o-V (0.248). The results highlight a consistent cross-lingual asymmetry: extracting answers from German documents is systematically harder when the query is in English (). However, the Agent mitigates this better in the open schema setting (0.384) compared to GPT-4o (0.292) and GPT-4o-V (0.212). The vision baseline struggles significantly here, likely because visual encoders often fail to parse the dense, multi-page tables typical of German DoPs.
3.3 Runtime, Cost and Token Comparison
Last, we compare the computational efficiency of the agentic system against both baseline models in terms of token usage, total cost, and runtime. Figure 3 summarizes these metrics across the KIE and QA tasks. Input and output token counts are derived from the prompt and completion segments. The cost is computed based on the GPT-4o pricing (September 2025), and the run time reflects the total wall clock time per tasks. Although AgenticIE provides performance benefits (see Section 3.1 and 3.2), it has the highest resource consumption. This is mainly due to the iterative nature of the agentic loop, where the planner and executor repeatedly exchange state information and tool outputs. High resource use of agentic AI is a known and open research problem Nooralahzadeh et al. (2024). Future work must focus on optimizing the planning horizon and reducing redundant tool calls to make agentic IE more sustainable for large-scale IE applications.
4 Related Work
Visually Rich Document Understanding. Layout-aware models Xu et al. (2020, 2021) and multimodal datasets Jaume et al. (2019); Park et al. (2019); Huang et al. (2019) combine text, layout, and visual features for structured extraction. However, many datasets and methods consist of clean, monolingual, and semantically regular documents. In contrast, DoP documents exhibit high variability in layout, quality, and language.
Prompt-Based Extraction. Prompt-based methods frame extraction as question answering Hu et al. (2023), incorporate layout structure Lin et al. (2023), and directly decode fields from images Kim et al. (2021). These methods reduce reliance on fixed schemas, but often assume consistently formatted, high-quality input. Further, they require a customization of the extraction pipeline to the specific task and dataset Colakoglu et al. (2025).
Tool-Augmented Reasoning and Modular Agents. Recent studies enhance LLMs with tool use Schick et al. (2023); Wang and Duan (2024), but most document pipelines remain static, relying on fixed OCR and parsing sequences Luo et al. (2024); Perot et al. (2023). More flexible designs Zhang et al. (2024a); Neupane et al. (2025) introduce partial modularity (e.g., knowledge guidance or layout-aware prompts), whereas they lack runtime planning or tool orchestration. Agentic systems Patil et al. (2023); Qin et al. (2023); Gao et al. (2023) decouple planning, tool use, and response generation. Benchmarks such as AgentBench Liu et al. (2023b) evaluate agents in interactive environments, but not for IE.
Planning, Feedback, and Memory in Agents. Recent patterns include task decomposition, plan revision, and tool routing Huang et al. (2024); Masterman et al. (2024). Iterative refinement methods implement feedback loops for recovery Shinn et al. (2023); Madaan et al. (2023); Liu et al. (2023a), which our agent mirrors via a planning-execution-feedback cycle to handle OCR or parsing failures. Structured planning frameworks support modular reasoning and collaborative workflows Wang et al. (2023); Wu et al. (2024), while code-trained LLMs offer stronger schema-aware planning Yang et al. (2024), motivating our choice of a structured planner. Long-term memory mechanisms, such as scratchpads or persistent state Zhang et al. (2024b), are integrated in our AgentState to track modality, task intent, and tool outcomes. Retrieval-augmented planners Ruan et al. (2023); Kong et al. (2023) also inform our approach to runtime tool selection across diverse document types.
5 Conclusion
This paper introduces AgenticIE, as the first agentic system and extensive dataset for IE from DoP documents, with a focus on handling hybrid schemas in English and German. Our agent outperforms static and multimodal LLM baselines, achieving higher EM scores compared to GPT-4o and GPT-4o-V across KIE and QA. The experimental analysis showcase the challenging nature of NLP applications to DoP.
Limitations
This work focuses on adaptive information extraction within a specific regulatory context-Declaration of Performance (DoP) documents-characterized by multilinguality and diverse layouts. The agentic system is therefore intentionally designed for domain-grounded reasoning rather than general-purpose intelligence. Extending it to other domains such as financial or legal records would require retraining the planner and adapting the tool registry to new schemas and reasoning conventions.
Our evaluation emphasizes structural and semantic correctness but does not address higher-order reasoning such as factual consistency, regulatory validation, or human-in-the-loop correction. Exploring such capabilities would shift the system from extraction toward auditable compliance reasoning. Likewise, while the agent performs reliably across English and German, scaling to low-resource or morphologically complex languages would demand cross-lingual schema alignment and translation-aware planning.
Finally, the current design focuses on textual content and inferred structure. Incorporating explicit visual cues such as layout topology or table geometry could further enhance precision in complex or scanned documents, but doing so would require multimodal model integration beyond this paper’s technical scope.
Ethical Considerations
The development and deployment of this agentic system for extracting information from Declaration of Performance (DoP) documents require careful consideration of several ethical dimensions, including data privacy, potential for misuse, algorithmic bias, and environmental impact.
Data Privacy and Personal Information.
As noted, DoP documents are public records mandated by EU regulation, but they can contain personal identifiers, such as the names, positions, and signatures of the individuals responsible for the declaration. While this information is publicly accessible, its automated collection and processing at scale warrant ethical scrutiny. Our approach mitigates privacy risks by using this data strictly for research purposes to evaluate the extraction system.
Reliability and Risk of Misuse
The information extracted from DoPs, such as fire resistance ratings or the use of hazardous substances, is critical for ensuring the safety and compliance of construction projects. An error in extraction could lead to the selection of non-compliant materials, posing significant safety risks. Therefore, this system is intended to be an assistive tool for engineers and architects, not a fully autonomous replacement for human oversight. End-users should be made aware that the system’s output must be verified by a qualified professional before being used for any critical decisions. The system’s built-in verification steps and high performance are designed to minimize errors, but the “human-in-the-loop” remains essential for accountability in high-stakes applications.
Algorithmic Bias and Fairness
Our dataset comprises documents in English and German. While the agentic system demonstrates more stable cross-lingual performance than the baselines, the paper notes that performance can still vary between languages and in cross-lingual scenarios. Deploying such a system across the multilingual European market could lead to performance disparities, potentially disadvantaging users working with less-represented languages. The data was also collected via web search and from specific repositories, which may introduce a selection bias towards certain manufacturers or product types. Future work should aim to expand the dataset to include a wider range of languages and document origins to ensure more equitable and robust performance.
Computational Cost and Environmental Impact
The experiments show that the proposed agentic system, while more accurate, incurs a significantly higher computational cost in terms of token usage, cost, and runtime compared to baseline models. The total token count for the agent was 38.69 million, costing over $300 and taking nearly 20 hours to run. We acknowledge that the energy consumption associated with such extensive computation has an environmental impact. This trade-off between performance and efficiency is a known challenge for agentic architectures. We advocate for continued research into more computationally efficient models and methods to create more sustainable solutions for information extraction.
AI Assistants In Research Or Writing
We have used GPT-5, Gemini 2.5 pro and Claude Opus 4.1 to help us refine our writing. All final text in the paper has been verified by the authors.
Appendix A Appendix
This appendix provides supplementary materials to support the reproducibility and deeper understanding of our AgenticIE system and the DoP dataset. We first detail the dataset construction process, including the target annotation keys, QA generation templates, and concrete examples of Key Information Extraction (KIE) and Question Answering (QA) pairs. Subsequently, we provide implementation specifics for the agentic architecture, covering the tool registry, prompt design strategies, and schema definitions. Finally, we present the full prompts used for the baseline models and a qualitative error analysis of extraction failures.
A.1 Dataset Annotation Details
A.1.1 Target Keys
The following JSON-style structure presents list of 12 target keys selected for annotation and extraction from DoP documents, including their English and German versions. These keys represent core elements required for regulatory compliance.
A.1.2 KIE Annotation Sample
Example of a multilingual KIE annotation for the key "Declared Performance" from a German-language DoP document.
The following JSON-style structure presents extracted key-value pairs (KIEs) from the document Sager_ISO-SWISS-Design-Panels.pdf, annotated in both English and German for evaluation and training purposes.
A.1.3 QA Dataset Generation Templates
The templates are organized into three categories: Hierarchical path-based templates (7 per language) leverage the full key path structure to generate context-aware questions where the question type (who, when, where) depends on the specific sub-field; Custom question mappings (2 per language) provide domain-specific phrasings for technical terminology such as AVCP systems; and Entity-type templates (2 per language) distinguish between person entities and generic attributes using "Who" versus "What" question forms. Template selection follows a hierarchical decision tree that identifies domain-specific patterns, checks for custom mappings, determines entity type, and applies fallback templates. Table 4 presents the complete English template set; German templates follow parallel structures.
| Type | Question Template |
| 1. Hierarchical Path-Based Templates (7) | |
| Sig-Name | Who signed the declaration ({key})? |
| Sig-Position | What is the signatory’s position ({key})? |
| Sig-Both | When and where was it signed ({key})? |
| Sig-Date | When was the declaration signed ({key})? |
| Sig-Place | Where was the declaration signed ({key})? |
| Decl-Perf | What is the value of ’{key}’? |
| Default | What is the {last_key} in ’{key}’? |
| 2. Custom Question Mappings (2) | |
| Tech-Doc | What technical documentation is provided? |
| AVCP | Which AVCP system applies? |
| 3. Entity-Type Templates (2) | |
| Person | Who is the {key}? |
| Generic | What is the {key}? |
Note: {key} represents the field name or path; {last_key} represents the final component of the hierarchical key path. German templates follow parallel structures (11 additional templates).
A.1.4 QA Annotation Sample
The following JSON-style structure presents a sample of question annotations derived from key-value pairs in the document 011-Leistungserklaerung-KS-10DF.pdf. Each entry specifies the key, its value, the associated question, and the language used.
A.1.5 Tool Registry and Modularization
To support robust and adaptable behavior across multilingual, layout-diverse, and partially scanned DoP documents, system relies on a modular tool architecture. Tools in this system represent discrete capabilities, such as PDF modality detection, OCR-based text extraction, question answering, or key-value pair retrieval, and are invoked selectively based on the agent’s current reasoning state.
At the heart of this architecture is a structured Tool abstraction, which wraps both deterministic utilities (e.g., OCR, PDF parsing) and LLM-powered capabilities (e.g., intent classification or verification) into a unified callable interface. Each tool object declares its name, functional implementation, Pydantic input schema, and optional pre- and post-processing logic. Tools also specify whether they require previous output, and which document types they are compatible with (e.g., scanned or text-based). This explicit declaration ensures that tools remain self-contained and declarative, while still integrating seamlessly into the planner-executor loop.
All tools are registered at startup by combining deterministic and LLM-based implementations into a central registry, forming a complete vocabulary of system capabilities. This registry remains fixed throughout execution and provides a filtered view of what actions are currently applicable, based on the evolving AgentState.
When the planner is invoked, it consults this registry to determine which tools are compatible with the current reasoning intent and document modality. Tools that require unavailable inputs (e.g., translation before text extraction) are excluded. The remaining candidates are embedded as tool metadata into the LLM prompt. The model then selects the most appropriate tool and input parameters via structured output, which is parsed and validated against the registry before dispatching it to the executor.
The executor interprets this plan by invoking the selected tool. Before execution, it applies a normalize_input() function to map the current state and raw parameters into a validated schema. Once executed, the result is post-processed via the tool’s postprocess() function, which updates the AgentState in a localized and semantically meaningful way. This decoupling of decision logic from execution logic ensures that tools remain modular, reusable, and interchangeable, while still preserving contextual awareness.
Unlike monolithic pipelines, the tools in the system are deeply integrated into a loop of stateful reasoning. For instance, modality detection tools such as check_if_scanned determine whether a document should be OCR’d or parsed directly, triggering either extract_text_ocr or extract_text_direct accordingly. Language inference via detect_language enables downstream translation and schema matching. Intent classification and target key parsing refine the agent’s task trajectory, while extraction tools like extract_key_values or answer_question carry out the core retrieval logic. The planner may later invoke verify_extraction to assess whether the output is complete and grounded in the source text, before optionally translating it into the user’s preferred language using translate_text. In edge cases, where noisy OCR output risks triggering LLM content filters, the system sanitizes text using sanitize_ocr_text as a recovery mechanism.
Each of these tools serves a specific role in a broader adaptive plan. Their activation is not hardcoded but emerges dynamically as the agent iterates through reasoning, acting, and revising. Tools modify the state incrementally, and their execution history is logged to support loop detection, misuse prevention, and multi-step recovery. Because each tool operates independently but updates a shared AgentState, the planner and executor can reason about partial progress, fallback strategies, or when to terminate altogether.
The tight integration of tools, registry, and reasoning enables the agent to dynamically adapt its behavior across a wide variety of document structures and user inputs, while preserving traceability and control. Figure 4 illustrates the complete tool orchestration pipeline, from initial tool registration to planner-driven selection and executor-based integration.
Multilingual QA and KIE Templates: To support multilingual document processing, the system employs language-specific prompt templates for both question answering and key-value extraction tasks. Based on the detected document_language, the agent selects between English (en) and German (de) versions of the prompts. These templates ensure that language constraints, extraction rules, and formatting guidelines are consistently enforced.
The examples below illustrate two representative prompts used during tool invocation: one for KIE extraction in English, and one for QA in German.
Each prompt includes a schema definition and task-specific extraction rules tailored to the document’s language. This modular design ensures that the model generates structured outputs consistent with downstream processing requirements such as verification and translation.
A.1.6 Prompt Design and Language Adaptation
The reasoning capability of our agentic system depends heavily on well-structured and context-aware prompts. These prompts guide both the planner and tool-specific language models to make decisions, extract structured information, and adapt across languages. This section outlines the design principles and implementation of prompt templates for planning, key-value extraction, question answering, error recovery, and multilingual handling.
Planner Prompt with State-Aware Reasoning: The planner prompt is dynamically constructed based on the current AgentState and tool metadata. The main structure is as follows:
It synthesizes the current AgentState into a rich context string that includes user intent, document metadata, extracted content, prior tool usage, tool descriptions, and reasoning history. This state-driven prompt is used to instruct a language model to determine the next action in the pipeline. The planner must always respond in a strict JSON format:
The prompt encodes decision-making rules directly, allowing for multi-step planning logic. For example, if the user intends to extract key-value pairs but the document is in German while the user expects English output, the planner prompt will contain structured rules for translation, extraction, verification, and result transformation:
Consider the state where the user has submitted a document titled DoP-1234.pdf with the intent of extracting key-values. The document has already been classified as scanned, and its language is detected as de (German), while the user’s preferred language is en (English). No extracted values or translated keys exist yet. The planner prompt will reflect these observations as extracted facts:
The planner might then respond with:
This strict format allows the system to introspect, log, and act deterministically on LLM outputs.
Error Recovery Prompts: To handle tool failures or malformed outputs, recovery prompts provide fallbacks. For instance, if the planner returns invalid JSON, the system uses get_malformed_response_prompt():
This prompt is structurally identical to the planner prompt but framed in an error-recovery mode, helping the agent self-correct without external intervention.
By combining language-specific prompt templates, strict JSON response schemas, and a planning prompt with embedded decision rules, the system achieves structured reasoning across multilingual documents. The modularity of this design also enables the easy addition of new languages, extraction schemas, or prompt variations, supporting scalability in diverse information extraction scenarios.
A.1.7 KIE Schema templates for EN/DE
The following templates define the expected JSON structure for the extraction task in German and English. They specify the hierarchy for complex fields like Declared Performance and Signature, ensuring consistent output formatting.
A.2 Baseline Prompt Examples
This section provides the full text of the prompts used for the baseline models (GPT-4o and GPT-4o-Vision). These prompts are designed to be strictly zero-shot, providing the model with the schema definition and formatting constraints.
A.2.1 KIE Baseline Prompt (Version 1 - English)
The English KIE prompt instructs the model to extract the 12 target keys into a flat or nested JSON structure, handling missing values with empty strings.
A.2.2 KIE Baseline Prompt (Version 2 - German)
The German KIE prompt is adapted to include the German target keys (e.g., Erklärte Leistung instead of Declared Performance) and specific language constraints to prevent anglicisms in the output.
A.2.3 QA Baseline Vision Prompt (English)
The Vision QA prompt is minimized to focus on visual reasoning, instructing the model to return concise string answers or lists without additional conversational filler.
A.3 Error Analysis
Table 5 provides a qualitative analysis of errors encountered during the Open Schema extraction task. As the system must dynamically infer nested key structures without a fixed template, the majority of failures stem from semantic mismatches rather than hallucinated values. Common error sources include: (1) Granularity Mismatches, where the agent predicts a parent key instead of a specific sub-field; (2) Omission of Qualifiers, such as missing units or specific descriptors (e.g., “Tabellenwert”); and (3) Spelling/OCR Errors in the final value. These examples illustrate the strict nature of the Exact Match metric, where even semantically equivalent schema predictions result in evaluation penalties.
| Ground Truth Key (GT) | Predicted Key | Explanation |
| Erklärte Leistung/Klasse der Brutto-Trockenrohdichte | Erklärte Leistung/Brutto-Trockenrohdichte | Qualifier “Klasse der” is missing in the prediction; keys fail to align. |
| Erklärte Leistung/Form und Ausbildung/Bezeichnung | Erklärte Leistung/Form und Ausbildung | Predicted key refers to a parent field; GT specifies a subfield. |
| Erklärte Leistung/Wasserdampfdurchlässigkeit (Tabellenwert) | Erklärte Leistung/Wasserdampfdurchlässigkeit | Suffix “(Tabellenwert)” is omitted; results in an exact match failure. |
| Erklärte Leistung/Maße/Länge [mm] | Erklärte Leistung/Maße | Prediction aggregates multiple subfields; GT expects specific ones. |
| Erklärte Leistung/mittlere Druckfestigkeit [N/mm²] | Erklärte Leistung/mittlere Druckfestigkeit | Unit annotation is missing in the prediction; leads to mismatch. |
| Unterschrift/Ort und Datum der Ausstellung | Unterschrift/Ort und Datum der Ausstellung | Keys match, but predicted value has a spelling error (“Lobnitz” vs. “Löbnitz”). |
| Erklärte Leistung/Wärmedurchlasswiderstand | — | Key missing in prediction; leads to evaluation exclusion. |
| Erklärte Leistung/Form und Ausbildung/Lochanteil von - bis [Vol %] | — | GT key has no corresponding prediction; reduces recall. |
| Erklärte Leistung/Grenzabmaße / Abmaßklasse | — | Missing predicted key causes evaluation penalty. |